OpenStudy (anonymous):

GUYS MEDALS AVAILABLE !!!! SOLVE PLS |-4x+24| = 2x-18

OpenStudy (campbell_st):

ummmm interesting solution... using the 1st equation x = 7 seems to be a solution.... but when you substitute it into the equation.... what happens...?

OpenStudy (anonymous):

rofl whoops, wrong way around, my bad

OpenStudy (anonymous):

okay, i got up to this point... what do i do now? : -(-4x+24) = 2x - 18

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

I put the negatives on the wrong side. They're meant to encapsulate the (2x-18)

OpenStudy (anonymous):

(-4x+24) = 2x - 18 (-4x+24) = -(2x - 18)

OpenStudy (campbell_st):

well if you distribute you get 4x - 24 = 2x - 18 subtract 2x from both sides 2x - 24 = -18 add 24 to both sides 2x = 6 solve for x then substitute it into the original equation to see if it makes sense

OpenStudy (campbell_st):

you should always check to see if the solution you got makes sense.... its one of the keys to success....

OpenStudy (anonymous):

THANKS GUYS, U GUYS R THE BEST <3

OpenStudy (nincompoop):

DOES IT MATTER IF YOU PUT THE NEGATIVE ON THE LEFT OR RIGHT HAND SIDE ?

OpenStudy (anonymous):

When I put the negative on the left hand side I got x = 3 & x = 7 on the right hand side, I got x=-1 & x=-21 none of which make sense

OpenStudy (nincompoop):

stick to the left hand

OpenStudy (nincompoop):

;)

OpenStudy (anonymous):

jeez